For the reaction H2(g) +Cl2(g) =2HCl(g) if 20 ml of H2 gas and 30ml of Cl2 gas are taken in a container then what is the volume of gas present in the container after the reaction

Grade:11

1 Answers

sam
6 years ago
I will do this by Mole concept as it will give conceptual clarity. 
we know that 1 mole of any gas occupies 22.4 L of volume at STP. 
in this reaction 1 millimole of Hgas is required to react with 1 millimole of Cl2 gas  to produce 2 millimoles of HCl gas
but since we do not have 1 millimole H2 , H2  will be the limiting regent.
Thus millimoles of H2 = 20/22.4
thus equal millimoles of Clare required = 20/22.4
therefore millimoles of HCl formed are 2\times 20/22.4 
thus volume of HCl formed = 2\times 20/22.4 \times 22.4 (as 1 mole of any gas takes 22.4 L of volume at STP)
But Clgas is also left in some quantity. That quantity is (30-20) ml = 10 ml
Volume of gas present in the container is 40 ml of HCl and 10 ml of CL
Thus total volume og gas is 50 ml